### v1

@@ 3.2 Parametric regularity for uniform random fields
LHS $\geq a_{\min}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}^{2}+c_{\min}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
$\geq\min(a_{\min},c_{\min})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2},\qquad c_{\min}:=\lambda-\kappa_{\max}>0.$ (3.12)
@@
For the right-hand side of (3.2), we pull out $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})$ and then apply the Cauchy–Schwarz inequality to the first two terms, as is typically done. The new challenge is to manage the space-time integrals of three factors involving $w$ in the last two terms of (3.2). If we blindly apply the Cauchy–Schwarz inequality multiple times to separate the three factors, then we would raise norms to higher powers than $2$ and the subsequent recursion would fail.
@@
To keep the norms with exponent $2$, we apply the generalized Hölder inequality, first for ${\boldsymbol{x}}\in\Omega$ with $\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1$, and then for $t\in I$ with $\frac{1}{\infty}+\frac{1}{2}+\frac{1}{2}=1$, as follows:
@@
$\leq\theta_{1}^{2}\,\bigg(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\bigg)\,\Big(\int_{I}\|g\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
@@
where we also used the embeddings (2.6) and (2.7). Hence we obtain from (3.2)–(3.2),
$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}$
$\leq\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\|_{L^{2}(I;L^{2}(\Omega))}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;L^{2}(\Omega))}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))},$
where we used $b\leq\kappa_{\max}\,e^{\lambda\,T}$. The last factor in every term on the right-hand side can be upper bounded by $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$. Canceling out this common factor from both sides and grouping the first two terms on the right-hand side, we arrive at
$\min(a_{\min},\lambda-\kappa_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
$\qquad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\,=:\,\Theta_{{\boldsymbol{\nu}}}.$ (3.14)
Time derivative. We return to (3.2) but remove the integration over $t\in I$. For a.e. $t\in I$ we have
@@
$\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}a\,\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v\,{\mathrm{d}}{\boldsymbol{x}}+\int_{\Omega}c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
@@
$=-\sum_{j\geq 1}\nu_{j}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}+\sum_{j\geq 1}\nu_{j}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}$
$\quad-\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\int_{\Omega}\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\$
$\quad-\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\int_{\Omega}\xi_{j}\,e^{\lambda\,t}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}.$
Using $0\leq w\leq e^{-\lambda\,t}$ to obtain $|2\,b\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,w\,v+c\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v|\leq|\kappa+\lambda+f|\,|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w|\,|v|$, and applying the Cauchy–Schwarz and Hölder inequalities, we obtain
$\Big|\int_{\Omega}\frac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\,v\,{\mathrm{d}}{\boldsymbol{x}}\Big|$
@@
$\leq a_{\max}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
@@
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
@@
$\quad+\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}$

### v2

@@ 3.2 Parametric regularity for uniform random fields
$\langle F_{\boldsymbol{\nu}},v\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}:=-\int_{\Omega}\bigg[\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v$ (3.8)
$+\kappa\,$ $\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,v+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\bigg]\,{\mathrm{d}}{\boldsymbol{x}},$
which will be shown to be well defined below. Since $u_{0}$ is independent of $\boldsymbol{y}$}, differentiating the initial condition in (2.12) gives the initial condition $\int_{\Omega}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0))\,v\,{\mathrm{d}}{\boldsymbol{x}}=0$ for all $v\in L^{2}(\Omega)$.
We now prove by induction that $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$ for all multiindices $\boldsymbol{\nu}$}. For the base case ${\boldsymbol{\nu}}={\boldsymbol{0}}$, we have $\partial_{\boldsymbol{y}}^{\boldsymbol{0}}u=u\in{\mathcal{X}}$ by Theorem 2.1. Then under the induction hypothesis that $\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\in{\mathcal{X}}$ for all ${\boldsymbol{m}}<{\boldsymbol{\nu}}$, it follows that $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega))^{*}=L^{2}(I;H^{1}(\Omega)^{*})$, which we show below by considering functionals corresponding to each term on the right-hand side of (3.8) separately.
Let $v\in L^{2}(I;H^{1}(\Omega))$. For the first term in (3.8) involving gradients, using $\psi_{j}\in L^{\infty}(\Omega)$ and $\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\in{\mathcal{X}}$ then applying the Cauchy–Schwarz inequality twice gives
$\int_{I}\int_{\Omega}\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\|\psi_{j}\|_{L^{\infty}(\Omega)}\int_{I}\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}\,{\mathrm{d}}t$
$\leq\|\psi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{{\mathcal{X}}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
Hence, the functionals corresponding to the gradient terms in (3.8) are in $L^{2}(I;H^{1}(\Omega)^{*})$. Similarly, for the second term in (3.8) we obtain
$\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\|\xi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{{\mathcal{X}}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
@@
For the third and fourth terms in (3.8), we face a new challenge to manage the space-time integrals of three factors involving $u$. If we blindly apply the Cauchy–Schwarz inequality multiple times to separate the three factors, then we would raise norms to higher powers than $2$ and the subsequent induction would fail.
@@
To handle integrals of the form $\int_{I}\int_{\Omega}f\,g\,h\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ where the integrand is a product of three factors and to keep the norms with exponent $2$, we apply the generalized Hölder inequality, first for ${\boldsymbol{x}}\in\Omega$ with $\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1$, and then for $t\in I$ with $\frac{1}{\infty}+\frac{1}{2}+\frac{1}{2}=1$, as follows:
@@
$\leq\theta_{1}^{2}\,\Big(\max_{t\in I}\|f\|_{L^{2}(\Omega)}\Big)\,\Big(\int_{I}\|g\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}\Big(\int_{I}\|h\|_{H^{1}(\Omega)}^{2}\,{\mathrm{d}}t\Big)^{1/2}$
@@
where we used the embeddings (2.7) and (2.8). Thus we obtain for the third and fourth terms in (3.8),
$\int_{I}\int_{\Omega}\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))},$
$\int_{I}\int_{\Omega}\xi_{j}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,v\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$ $\leq\theta_{1}^{2}\,\theta_{2}\,\|\xi_{j}\|_{L^{\infty}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|v\|_{L^{2}(I;H^{1}(\Omega))}.$
Seeing as though $F_{\boldsymbol{\nu}}$ is a linear combination of functionals belonging to $L^{2}(I;H^{1}(\Omega)^{*})$, we conclude that $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega)^{*})$.
Since the bilinear form corresponding to the second term in (3.7) is elliptic and bounded and $F_{\boldsymbol{\nu}}\in L^{2}(I;H^{1}(\Omega)^{*})$, it then follows from standard results for linear parabolic PDEs (e.g., [12, 45]) that there exists a unique $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$ satisfying (3.7).
Gradient. Taking the mixed derivative $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}$ on both sides of the second condition in (2.12) and then taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)$, we conclude that $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$. Since $\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\in{\mathcal{X}}$, by (2.6), this yields
@@
$\int_{0}^{\tau}\!\!\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\,{\mathrm{d}}t=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}{\mathrm{d}}t=\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}}{2}\geq 0.$
Taking $v=\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u$ in (3.6) then gives
@@
$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}{\mathrm{d}}t+\!\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}\,u+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}+\eta\Big]{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t=0,$ (3.10)
where
$\eta({\boldsymbol{x}},t)$ $:=\sum_{j\geq 1}\nu_{j}\,\psi_{j}\,\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\cdot\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)-\sum_{j\geq 1}\nu_{j}\,\xi_{j}\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)$ (3.11)
$\quad+\kappa\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)$
$\quad+\sum_{j\geq 1}\nu_{j}\,\xi_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,(\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u)\,(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u).$
Note that $\eta$ can potentially give both positive and negative values.
Until this point we still have equality in (3.10)–(3.11). For the special cases $|{\boldsymbol{\nu}}|=1$, the sum over ${\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}$ is empty and takes the value 0, with no meaning attached to the summand.
Recall that $0\leq u\leq 1$. Now we drop the nonnegative terms $\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}\,u$ and $f\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)^{2}$ in (3.10) and rearrange to obtain the inequality
@@
$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+\eta\Big]\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\kappa_{\max}\int_{0}^{\tau}\!\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(\Omega)}^{2}\,{\mathrm{d}}t.$
If we were to return to the derivation of (2.12), but multiply the PDE by $e^{-2\,\kappa_{\max}\,t}$ before integrating over $t\in[0,\tau]$, then the preceding argument would lead to the above inequality but with all integrands multiplied by $e^{-2\,\kappa_{\max}\,t}$, and by the product rule it is equivalent to
@@
$\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\Big(\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,t}\Big)\,{\mathrm{d}}t+\int_{0}^{\tau}\!\!\int_{\Omega}\Big[a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}+\eta\Big]e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq 0.$ (3.12)
Using $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,0)\|_{L^{2}(\Omega)}^{2}=0$ and (2.6), we drop the nonnegative term $a\,|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)|^{2}$ in (3.12) to obtain
$\frac{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,\tau)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,\tau}}{2}=\int_{0}^{\tau}\!\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\Big(\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}^{2}e^{-2\,\kappa_{\max}\,t}\Big)\,{\mathrm{d}}t\leq\int_{0}^{\tau}\!\!\int_{\Omega}|\eta|\,e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$
We multiply both sides of the inequality by $2\,e^{2\,\kappa_{\max}\,\tau}$ and integrate over $\tau\in I$ to obtain
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}^{2}$ $\leq 2\int_{0}^{T}\!\!\int_{0}^{\tau}\!\!\int_{\Omega}|\eta|\,e^{2\,\kappa_{\max}\,(\tau-t)}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\,{\mathrm{d}}\tau=\int_{I}\int_{\Omega}|\eta|\,\frac{e^{2\,\kappa_{\max}\,(T-t)}-1}{\kappa_{\max}}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.13)
Next we drop the nonnegative first term in (3.12) and take $\tau=T$ to obtain
$a_{\min}\,e^{-2\,\kappa_{\max}\,T}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}^{2}\leq\int_{I}\int_{\Omega}|\eta|\,e^{-2\,\kappa_{\max}\,t}\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.14)
Combining (3.13) and (3.14), we conclude that
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}=\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}^{2}+\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}^{2}$
$\leq\int_{I}\int_{\Omega}|\eta|\,\Big(\frac{e^{2\,\kappa_{\max}\,(T-t)}-1}{\kappa_{\max}}+\frac{e^{2\,\kappa_{\max}\,(T-t)}}{a_{\min}}\Big)\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t\leq\frac{2\,e^{2\,\kappa_{\max}\,T}}{\min(a_{\min},\kappa_{\max})}\int_{I}\int_{\Omega}|\eta|\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t.$ (3.15)
To bound the space-time integral of $|\eta|$ with $\eta$ defined in (3.11), we pull out $\beta_{j}:=\max(\|\psi_{j}\|_{\infty},\|\xi_{j}\|_{\infty})$ and then apply the Cauchy–Schwarz inequality to the first two terms, as is typically done. For the last two terms in (3.11) we again use (3.2) to manage the space-time integrals of three factors. We obtain from (3.11), (3.2), and (3.2),
$\frac{\min(a_{\min},\kappa_{\max})}{2\,e^{2\,\kappa_{\max}\,T}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}\leq\int_{I}\int_{\Omega}|\eta|\,{\mathrm{d}}{\boldsymbol{x}}\,{\mathrm{d}}t$
$\leq\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)\|_{L^{2}(I;L^{2}(\Omega))}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{L^{2}(I;L^{2}(\Omega))}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;L^{2}(\Omega))}$
@@
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$